A total of 825 tickets were sold for the school play they were either adult tickets or student tickets they were 75 much do tickets sold in adult tickets how many adult tickets were sold


1
Expert's answer
2021-12-13T16:43:14-0500

"\begin{cases}a+s=825\\a-s=75 \end{cases}"

"2a=900\to a=\frac{900}{2}=450."

"s=a-75=450-75=375."

450 adult tickets were sold.